# Leveraging Edge Intelligence and LLMs to Advance 6G-Enabled Internet of Automated Defense Vehicles

### v1

@@ V-D Cybersecurity Concerns
In another scenario, consider a military AV equipped with an LLM for navigation and communication. If an attacker were to employ a prompt injection attack, the AV could be misled into executing commands that divert it into ambush zones or cause it to ignore critical threats. Such a breach could compromise mission integrity, endanger personnel, and reveal operational details to adversaries. Connected vehicles in the network can suffer from various malicious attacks that are used for either information leakage or operation sabotage. For example, ADVs rely on an array of sensors such as cameras, LiDAR, RADAR, and so on to perceive their environment, and these sensors are vulnerable to various forms of manipulation and attacks, such as signal jamming and sensor spoofing. Moreover, fake updates or Interruptions during critical updates by malicious actors prevent ADVs from performing proper actions and harm the operation.
@@ V-E LLM Inherited-Limitations
One major risk of using LLMs in military applications is ”hallucination,” where the model produces outputs that are factually incorrect or entirely fabricated. This poses a serious threat in high-stakes battlefield situations, where decisions based on false information can have catastrophic results. In dynamic combat scenarios, if commanders depend on LLMs for real-time analysis, a hallucination about crucial details—such as enemy positions or friendly force status—could result in misguided decisions, leading to missed opportunities or unnecessary engagements.
@@ V-F Cost of Deployment and Integration of 6G
The deployment and integration of 6G and MEC for ADVs come with significant costs, which include infrastructure upgrades, such as installing 6G base stations, building edge data centers, and upgrading backhaul networks. ADVs also need advanced hardware, such as 6G-compatible modems, antennas, and edge computing modules. These systems along with high-end sensors and AI-driven software can increase vehicle costs. Operational costs such as network maintenance, energy consumption, and secure data transfer will further add to the overall investment. Additionally, military operations are performed not only in urban areas but also in primitive areas where there are fewer technological infrastructures in dangerous zones. ADVs should also handle their tasks in these situations.

### v2

@@ V-D Cybersecurity Concerns
In another scenario, consider a military AV equipped with an LLM for navigation and communication. If an attacker were to employ a prompt injection attack, the AV could be misled into executing commands that divert it into ambush zones or cause it to ignore critical threats. Such a breach could compromise mission integrity, endanger personnel, and reveal operational details to adversaries. Connected vehicles in the network can suffer from various malicious attacks that are used for either information leakage or operation sabotage [5]. For example, ADVs rely on an array of sensors such as cameras, LiDAR, RADAR, and so on to perceive their environment, and these sensors are vulnerable to various forms of manipulation and attacks, such as signal jamming and sensor spoofing. Moreover, fake updates or Interruptions during critical updates by malicious actors prevent ADVs from performing proper actions and harm the operation. Therefore, Secure communication methodologies should utilize military-grade encryption like AES-256 and quantum-resistant cryptographic techniques to prevent cyber threats and ensure mission-critical data integrity.
@@ V-E LLM Inherited-Limitations
One major risk of using LLMs in military applications is ”hallucination,” where the model produces outputs that are factually incorrect or entirely fabricated. This poses a serious threat in high-stakes battlefield situations, where decisions based on false information can have catastrophic results. In dynamic combat scenarios, if commanders depend on LLMs for real-time analysis, a hallucination about crucial details—such as enemy positions or friendly force status—could result in misguided decisions, leading to missed opportunities or unnecessary engagements. Mitigating hallucination risks in LLM models ensures the reliability and accuracy of their outputs and it can be provided by the several model adaptation techniques. Retrieval-augmented generation (RAG), which enhances models with external knowledge sources, ensures that AI outputs are grounded in factual data. Moreover, fine-tuning domain-specific data and adjusting pre-trained models with high-quality, domain-specific datasets aligns the LLM’s responses with verified and contextually accurate knowledge. Automated Reasoning, which is a verification mechanism, and employing mathematical proofs to validate AI system outputs can enhance trustworthiness. [7]
@@ V-F Cost of Deployment and Integration of 6G
The deployment and integration of 6G and MEC for ADVs come with significant costs, which include infrastructure upgrades, such as installing 6G base stations, building edge data centers, and upgrading backhaul networks. ADVs also need advanced and rugged hardware including reinforced chassis, temperature-resistant electronics, and shock-proof sensors, to withstand extreme operational environments and compatible with advance communications, such as 6G-compatible modems, antennas, and edge computing modules. These systems along with high-end sensors and AI-driven software can increase vehicle costs. Operational costs such as network maintenance, energy consumption, and secure data transfer will further add to the overall investment. Additionally, military operations are performed not only in urban areas but also in primitive areas where there are fewer technological infrastructures in dangerous zones or environments where infrastructures failure due to hostile actions.
Though terahertz (THz) communications adaptation can provide reliability by providing uninterrupted connectivity, THz signals are highly susceptible to environmental factors such as atmospheric absorption, precipitation, and obstacles, which can significantly attenuate the signal strength, especially in harsh environments, including remote or combat zones. To mitigate these issues, adaptive communication strategies, such as real-time environmental sensing and dynamic beamforming, can be employed to intelligently adjust transmission parameters in response to changing conditions. On the other hand, in the case of infrastructure failure, fallback mechanisms such as multi-layer network redundancy, Satellite or high-altitude platform (HAP)-based backup communication, mobile ad-hoc networks (MANETs) and device-to-device (D2D) communication are practical solutions to maintain operational continuity, providing alternative sources and backup of the service for critical conditions [5].